find the roots using newton rhapson. Stop until at 15th iteration or when approximate percent relative error is below 0.05%.

2x^5+x^4-2x-1=0
1
Expert's answer
2020-10-22T17:50:28-0400
"2x^5+x^4-2x-1=0""x^4(2x+1)-(2x+1)=0""(2x+1)(x^4-1)=0""(2x+1)(x^2+1)(x^2-1)=0""(2x+1)(x^2+1)(x+1)(x-1)=0"

"x1=-1, x2={1\over 2}, x3=1"

"f(x)=2x^5+x^4-2x-1"


"f'(x)=10x^4+4x^3-2"

"x_{n+1}=x_n-\dfrac{f(x_n)}{f'(x_n)}"

Initial solution "x_0=0"


"f(0)=-1, f'(0)=-2, x_1=0-{-1 \over -2}=-0.5""\begin{matrix}
 n & x_n & f(x_n) & f'(x_n) & x_{n+1} & f(x_{n+1}) \\
 1 & -0.5 & 0 & -1.875
\end{matrix}"

"x=-0.5"


Initial solution "x_0=2"


"f(2)=75, f'(0)=190, x_1=2-{75 \over190}=1.60526""\begin{matrix}
 n & x_n & f(x_n) & x_{n+1} & f(x_{n+1}) \\
 1 & 1.60526 & 23.74820 & 1.31189 & 7.10986 \\
 2 & 1.31189 & 7.10986 & 1.11790 & 1.81772 \\
 3 & 1.11790 & 1.81772 & 1.02326 & 0.29343 \\
 4 & 1.02326 & 0.29343 & 1.00111 & 0.01332 \\
 5 & 1.00111 & 0.01332 & 1.00000 & 0.00003\\
 6 & 1.00000 & 0 & 1.00000 
\end{matrix}"



"\varepsilon =\big|\dfrac{n_{i+1}-n_i}{n_i}\big|\cdot 100\%"

"\varepsilon =\big|\dfrac{1.31189-1.60526}{1.60526}\big|\cdot 100\%=18.26\%"

"\varepsilon =\big|\dfrac{1.11790-1.31189}{1.31189}\big|\cdot 100\%=14.79\%"

"\varepsilon =\big|\dfrac{1.02326-1.11790}{1.11790}\big|\cdot 100\%=8.47\%"

"\varepsilon =\big|\dfrac{1.00111-1.02326}{1.02326}\big|\cdot 100\%=2.16\%"

"\varepsilon =\big|\dfrac{1.00000-1.00111}{1.00111}\big|\cdot 100\%=0.11\%"

"\varepsilon =\big|\dfrac{1.00000-1.00000}{1.00000}\big|\cdot 100\%=0\%"

"x=1"